OpenStudy (anonymous):

Three times the smallest of three consecutive odd integers is the same as 1 subtracted from twice the largest number. Find all three integers.

OpenStudy (anonymous):

Let the three consecutive odd integers be : x-1, x, x + 1 So. According to the question: \(3 \cdot (x - 1) = 2 \cdot (x + 1) - 1 \) \(3x - 3 = 2x + 2 - 1\) \(3x - 3 = 2x + 1\) Can you solve this for x ??

OpenStudy (anonymous):

How would the x's all be odd if you do it like this?

OpenStudy (anonymous):

Oh what the hell... Wait..

OpenStudy (anonymous):

Consecutive odd integers will be : x, x + 2, x + 4

OpenStudy (anonymous):

lol good question. Why didn't you make your numbers to be x, x+2 and x+4

OpenStudy (anonymous):

Or you can let them as : x - 2, x , x + 2...

OpenStudy (anonymous):

But when you solve for the x in the last equation, x=4

OpenStudy (anonymous):

That is wrong @katoney55

OpenStudy (anonymous):

So the equation now would be 3 X (x+2) = 2 X (x+4) - 1

OpenStudy (anonymous):

Let the three consecutive odd integers be : x - 2, x and x + 2. So: \(3 \cdot(x - 2) = 2 \cdot (x + 2) - 1\) \(3x - 6 = 2x + 3\) Solve this..

OpenStudy (anonymous):

Yes the equation will be that now.. Solve that @katoney55

OpenStudy (anonymous):

Or let me try using x, x + 2 and x + 4 \[3x = 2(x + 4) -1 \implies 3x = 2x + 7 \implies x = 7\]

OpenStudy (anonymous):

Awesome! Thank you so much @waterineyes

OpenStudy (anonymous):

1, 3 & 5 are your odd numbers

OpenStudy (anonymous):

Solve that @katoney55 and check if you are getting x as 9..

OpenStudy (anonymous):

*5

OpenStudy (anonymous):

I put in x, x+2, and x-2 I got -1, -3, and -5.

OpenStudy (anonymous):

Order is important..

OpenStudy (anonymous):

Smallest number will be : (x - 2) and largest number will be : ( x + 2)..

OpenStudy (anonymous):

Can you show how you got -1 -3 and -5 ??

OpenStudy (anonymous):

3 X (x-2) +2 X (x+2) -1

OpenStudy (anonymous):

\[3 \cdot(x - 2) = 2 \cdot (x + 2) - 1\] \[3x - 6 = 2x + 3\] Solve this for x..

OpenStudy (anonymous):

\[3x - 2x = 6 + 3 \implies x = 9\]

OpenStudy (anonymous):

So the numbers will be : \[9 -2, 9, 9+2\]

OpenStudy (anonymous):

\[3(2x+1)=2(2x+5)-1\]Solve for x. x=3. Set x = 3 in the following list elements and evaluate: {(2 x + 1), (2 x + 3), (2 x + 5)} {7, 9, 11}

OpenStudy (anonymous):

@mayayamfundo three times the smallest number means 3x..

OpenStudy (anonymous):

7, 9 ,11. I see what I did wrong, I subtracted from one side when I was suppose to add!

OpenStudy (anonymous):

7, 9,11! I got it! thank you!

OpenStudy (anonymous):

You are adding the three numbers but that is not what question is saying: it is saying that: 3x = 2(x + 4) - 1

OpenStudy (anonymous):

oooops, my bad. Misreading

OpenStudy (anonymous):

Yes well done @katoney55

OpenStudy (anonymous):

It is okay...

OpenStudy (anonymous):

Thank you @waterineyes

OpenStudy (anonymous):

Welcome dear.. @robtobey you were doing right earlier why you deleted that ??

OpenStudy (anonymous):

Three times the smallest of three consecutive odd integers is the same as 1 subtracted from twice the largest number. 3*7 = 11*2 - 1 21 = 21

OpenStudy (anonymous):

Yep verification is right...

OpenStudy (anonymous):

@waterineyes I was hurrying along and was not sure of the calculations.

OpenStudy (anonymous):

Okay..